OpenStudy (anonymous):

a square poster has sides measuring 2 ft. less than the sides of a square sign. if the difference between their ares is 44 sq. ft. find the lengths of the sides of the poster and the sign

OpenStudy (anonymous):

p = s - 2, s^2-p^2 = 44

OpenStudy (anonymous):

That gives you 44 = s^2 - (s-2)^2 = 4s - 4. Add 4 and divide by 4 to give s = 12. So the sign has side 12 and the poster has side 10.

OpenStudy (anonymous):

thnk u